OpenStudy (anonymous):

A marble is drawn at random from the bag shown below. The bag contains 3 green, 4 yellow, 5 blue, and 8 pink marbles. What is P(not pink)?

jimthompson5910 (jim_thompson5910):

The pic isn't needed really since you are given "The bag contains 3 green, 4 yellow, 5 blue, and 8 pink marbles."

jimthompson5910 (jim_thompson5910):

how many marbles are there total?

OpenStudy (anonymous):

19 marbles in all

jimthompson5910 (jim_thompson5910):

close but no

OpenStudy (anonymous):

20

jimthompson5910 (jim_thompson5910):

how many are pink?

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

so 20 - 8 = 12 are not pink

jimthompson5910 (jim_thompson5910):

making P(not pink) = (# of non pink marbles)/(# of marbles total) P(not pink) = 12/20 P(not pink) = 3/5 ... make sure to reduce fully

OpenStudy (anonymous):

OMGG THANK YOU SO MUCHH I GET IT

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking now

OpenStudy (anonymous):

Lol